# Welfare and Cost Aggregation for Multi-Agent Control: When to Choose Which Social Cost Function, and Why?

### v2

@@ Abstract
Many multi-agent socio-technical systems rely on aggregating heterogeneous agents’ costs into a social cost function (SCF) to coordinate resource allocation in domains like energy grids, water allocation, or traffic management. The choice of SCF often entails implicit assumptions and may lead to undesirable outcomes if not rigorously justified. In this paper, we demonstrate that what determines which SCF ought to be used is the degree to which individual costs can be compared across agents and which axioms the aggregation shall fulfill. Drawing on the results from social choice theory, we provide guidance on how this process can be used in control applications. We demonstrate which assumptions about interpersonal utility comparability – ranging from ordinal level comparability to full cardinal comparability – together with a choice of desirable axioms, inform the selection of a correct SCF, be it the classical utilitarian sum, the Nash SCF, or maximin. We then demonstrate how the proposed framework can be applied for principled allocations of water and transportation resources.
I Introduction
@@
Multi-agent socio-technical control applications with heterogeneous agents arise in various domains, such as energy grids, traffic control, water distribution, and bandwidth allocation [1]. In these systems, the control objective in the form of a social cost function (SCF) is typically context-specific, and it often involves some cost or utility aggregation across agents, reflecting agents’ different goals, needs, and operational constraints. When performance indicators (e.g., energy consumption, travel time, bandwidth usage) are measurable and objective, one might sometimes quite straightforwardly design an SCF such as total cost or throughput. However, when agents’ costs represent subjective, agent-specific valuations, it is often unclear if these costs can be compared objectively, leading to difficulties and ambiguities in aggregating these costs and defining SCFs.
@@
Several candidate SCFs, including the sum (or weighted sum) of utilities/costs, Nash social cost, and max-min objectives, exist in the literature [2, 3, 4, 5], with the choice among these criteria usually depending on desired properties such as tractability, fairness, and robustness. A most popular approach to aggregate individual costs is the classical utilitarian rule, which sums individual costs into a single objective. While intuitive and computationally tractable, this approach assumes that the cost for one agent is commensurate with the cost of another. In practice, this is a strong comparability assumption which can lead to unintended consequences [6], e.g., disproportionally high wait times for ride-hails in remote areas [7], discrimination against certain train types or routes in real-time train re-scheduling [8], or exacerbation of energy poverty [9].
Exact full comparability of the costs incurred by different agents can be impeded by various reasons. In some cases, it is an issue of measurability, i.e., it might be difficult to precisely estimate costs because they depend on complex models and unknown parameters (e.g., the yield of a farm as a function of the water allocated to it or the delay incurred outside the traffic network of interest). In some cases, the decision maker cannot compare the subjective evaluation of the costs incurred by an agent (e.g., the value of travel time, which depends on socio-economic and circumstantial factors[10, 11]). In other cases, costs are self-reported and therefore vulnerable to strategic manipulation (e.g., customers declaring their electricity needs [12]). Finally, limited comparability of agents’ costs may encode fairness criteria, as it defines what information about the agents’ preferences should be considered in the decision (e.g., a deliberate decision on whether a traffic congestion protocol should prioritize vehicles that have accumulated delay).
@@
    Determine the level of interpersonal comparability. Such a decision is guided by the measurability and comparability of the agents’ preferences but it can also be a deliberate decision guided by politics, social agreements, and public sentiment.
@@
    Select the appropriate social cost function. Based on the selected comparability level, a restricted choice of possible SCFs is allowed. We provide rigorous guarantees about its robustness with respect to the lack of measurability and comparability.
@@
    Apply social cost function. Given an appropriate control, decision or optimization problem, the selected SCF can be explicitly optimized to select the best social choice, or can be used to for the analysis and design of these engineering systems. We provide rigorous guarantees about the set of permissible operations.
@@
We expect these results to be relevant in many control applications where specifications are provided in the form of social cost/welfare function [13]. The mathematical guarantees that we provide can be incorporated in the design of system-wide metrics that are now being used in multi-agent control, like Price of Anarchy [14, 15, 16, 17, 18]. They can also guide the design of optimization dynamics [19, 20, 21] and resource allocation mechanisms [22, 23]. Finally, our work responds to the growing interest in fairness in control [1, 24, 25, 26, 27, 28, 29], as many fairness notions arise naturally from careful considerations of the level of interpersonal comparability [29]. Indeed, this approach can also clarify the implicit fairness assumptions in existing game-theoretic solution concepts, such as the variational Generalized Nash Equilibrium, which has been shown to rely on a strong, and often unstated, comparability assumptions that, if violated, can lead to undesirable ’fairness’ outcomes [30].
@@
The aforementioned three-step procedure can also be interpreted in reverse, especially when designing engineering systems: if a specific notion of efficiency/fairness is desired (utilitarian, min/max, etc.) then a sufficient level of comparability needs to be achieved, which may require the collection of additional information from the agents, the design of a manipulation-safe mechanism, or the use of more accurate cost models. This may give a principled perspective on how freely agents’ costs can be designed and engineered [31, 32, 33].
The remainder of this paper is structured as follows. In Section II, we introduce desirable axioms that will set the stage for a principled welfarist approach based on social choice theory [34, 35, 36]. In Section III, we review and extend results on how desirable axioms, together with limits/possibilities of interpersonal cost comparability, determine which social cost aggregations and operations on them are permissible. In Section IV, we illustrate the versatility of the proposed framework via examples from water distribution, transportation, and energy.
II Welfarism
@@
In this section we introduce preliminaries adopted from social choice theory that allow a rigorous aggregation of individual costs and derivation of social cost functions. We lay out the foundations of a so called “welfarist´’ approach, that requires that all the relevant information for the social decision is contained in the agents’ cost functions.
II-A Preliminaries and Axiomatic Foundations
@@
Let $\mathcal{N}=\{1,2,\dots,n\}$ be a finite set of agents and let $\mathcal{X}$ denote the set of feasible outcomes $x$. An outcome can be a specific allocation of a scarce resource or any control decision that affects the agents. Each agent $i$ attaches a (possibly negative) cost $J_{i}:\mathcal{X}\to\mathbb{R}$ to every outcome, lower numbers being preferable, with $J_{i}(\cdot)\in\mathcal{J}$, where $\mathcal{J}$ is the set of all real-valued functions on $\mathcal{X}$. We let $\mathbf{J}=(J_{1},\dots,J_{n})$ denote a profile of such costs for the entire set of agents. When formulating a resource allocation problem that depends on agents’ individual evaluations represented by $\mathbf{J}$, one needs to define a single social preference relation $\succsim_{\mathbf{J}}$ on $\mathcal{X}$, so as to capture the collective or “social” viewpoint.
@@
Before determining how to incorporate each agent’s evaluation into a collective decision, one should first specify which fundamental properties (or axioms) this social preference relation $\succsim_{\mathbf{J}}$ needs to satisfy. Two classical well-established properties that can be expected from such a relation are:
@@
In control applications, it is convenient to work with real-valued measures of social cost rather than a social ordering of outcomes. This requires an additional mild continuity assumption, which postulates that if an outcome $x$ is socially strictly preferred over $y$, it should remain so under a small enough perturbation of the individual costs. As we show in Section III-B, this condition is easily verified for the setting considered in this work.
Definition 1 (Pairwise Continuity (PC)).
For every $\varepsilon\in\mathbb{R}_{++}^{N}$, there exists $\varepsilon^{\prime}\in\mathbb{R}_{++}^{N}$ such that for every profile $\mathbf{J}$ and every pair $x,y\in\mathcal{X}$ with $x\,\succ_{\mathbf{J}}\,y$, there exists a profile $\tilde{\mathbf{J}}$ satisfying $\tilde{\mathbf{J}}(x)\geq\mathbf{J}(x)+\varepsilon^{\prime}$ and $\tilde{\mathbf{J}}(y)\leq\mathbf{J}(y)+\varepsilon$, and such that $x\,\succ_{\tilde{\mathbf{J}}}\,y.$
In some cases it may be desirable to enforce an additional axiom that does not allow to distinguish individual agents.
@@
II-B Welfarism
If we accept these axioms (P) and (IIA) as reasonable and impose them on our social preference relation, together with condition (PC), we can establish a welfarist principle [37, 34], that implies that any complete and transitive social ordering can be expressed through a Social Cost Function (SCF) $C$, which subsumes all relevant information for the ordering.
Lemma 1 (Welfarism; Thm 3.7, [36], Thm 1, [35]).
@@ II-B Welfarism
The SCF $C$ thus represents the SCFL $\mathfrak{F}$ and all relevant information for ranking outcomes is contained in the $n$-tuple $\bigl(J_{1}(x),\dots,J_{n}(x)\bigr)$. The welfarist approach thus states that once $\mathbf{J}$ and the feasible set $\mathcal{X}$ are specified, the social preference depends only on the vector of costs for each outcome. Any other features of $x$ or $y$ are irrelevant from the perspective of the social ordering.
@@
As mentioned in the introduction, several well-known examples of SCFs $C$ have been proposed to aggregate individual costs. For instance, the utilitarian approach, often attributed to Bentham [38] as later formalized by Harsanyi [39], aggregates costs via a sum: $C_{\mathrm{util}}(J_{1},\dots,J_{n})=\sum_{i=1}^{n}J_{i}$. By contrast, the Rawlsian or min-max rule, inspired by Rawls [40], focuses on the well-being of the worst-off agent: $C_{\mathrm{Rawls}}=\max_{i\in\mathcal{N}}J_{i}$. Another example is the Nash Social Welfare function [41, 42], which is defined as $C_{\mathrm{Nash}}(u_{1},\dots,u_{n})=-\prod_{i=1}^{n}(-J_{i})$. Notably, these three canonical SCFs can all be viewed as members of the family of Hölder means. Generally, a large body of literature guides the selection of a SCF based on the desirable properties reflecting notions of fairness, equity (e.g. envy-freeness and its relaxations, proportional fairness, Pigou-Dalton principle etc.), efficiency, or some trade-off between the two, reflected in measures such as maximin share (MMS) guarantee, alpha-fairness, or statistical inequality indices (e.g., the Gini coefficient, coefficient of variation, or Hoover index) [43, 5, 6, 44]. Instead, the focus of this work is to provide a “first-principle” way to make this choice by first considering the fundamental assumptions of interpersonal comparability of the agents’ costs, which are needed to properly define “efficiency” and “fairness” in the first place, and from which the appropriate aggregation rule naturally follows.
III Cost Aggregation and Comparability
@@
The welfarist principle (Lemma 1) tells us that any social ordering under (P), (IIA) and (PC) must be representable by a SCF that depends solely on the costs incurred by the agents. One must still decide which $C$ is appropriate in practice. Such a choice needs to be guided by a deliberate decision on how individual costs can be measured and compared across agents, so that meaningful interpersonal trade-offs can be made in the social choice. Addressing the interpersonal comparability issue corresponds to deciding which transformations of the individual costs leave the social ranking unchanged [34, 35]. By restricting or expanding the class of invariance transformations of the individual costs, one effectively selects a level of measurability and comparability for the agents’ costs. Once the comparability assumptions are fixed, the shape of the SCF $C$ is essentially pinned down.
This section formalizes these invariance principles and shows how each comparability assumption leads to distinct classes of SCFs.
III-A Comparability Levels
Definition 2 (Invariance Transformations under a SCF).

### v3

# Position Paper

@@ Abstract
Many multi-agent socio-technical systems rely on aggregating heterogeneous agents’ costs into a social cost function (SCF) to coordinate resource allocation in domains like energy grids, water allocation, or traffic management. The choice of SCF often entails implicit assumptions and may lead to undesirable outcomes if not rigorously justified. In this paper, we demonstrate that what determines which SCF ought to be used is the degree to which individual costs can be compared across agents and which axioms the aggregation shall fulfill. Drawing on the results from social choice theory, we provide guidance on how this process can be used in control applications. We demonstrate which assumptions about interpersonal utility comparability - ranging from ordinal level comparability to full cardinal comparability - together with a choice of desirable axioms, inform the selection of a correct SCF, be it the classical utilitarian sum, the Nash SCF, or maximin. Thus, fixing comparability level first, then choosing an objective from the compatible class, and reporting both as part of the specification, makes the fairness and efficiency consequences transparent. We demonstrate how the proposed framework can be applied for principled allocations of water, transportation, and energy resources.
{IEEEkeywords}
multi-agent control, social cost function, resource allocation, utilitarian, Nash, maximin, interpersonal comparability
1 Introduction
@@
Multi-agent socio-technical control applications with heterogeneous agents arise in various domains, such as energy grids, traffic control, water distribution, and bandwidth allocation [CSSroadmap]. In these systems, the control objective in the form of a social cost function (SCF) is typically context-specific, and it often involves some cost or utility aggregation across agents, reflecting agents’ different goals, needs, and operational constraints. When performance indicators (e.g., energy consumption, travel time, bandwidth usage) are measurable and objective, one might quite straightforwardly design an SCF such as total cost or throughput. However, when agents’ costs represent subjective, agent-specific valuations (e.g. electricity needs of a hospital and a household, value of time for a population of travelers with heterogeneous income etc.), it is often unclear if these costs can be compared on a common scale, leading to difficulties and ambiguities in aggregating these costs and defining SCFs.
@@
Several candidate SCFs, including the sum (or weighted sum) of costs, Nash social cost, and min-max objectives, exist in the literature [caragiannis2019mnw, ramezani_endriss_2010, radunovic2007unified, bertimas2011priceoffairness], with the choice among these criteria usually depending on desired properties such as tractability, fairness, and robustness. The most popular approach to aggregate individual costs is the classical utilitarian rule, which sums individual costs into a single objective. While intuitive and computationally tractable, this approach assumes that the cost for one agent is commensurate with the cost of another. In practice, this is a strong comparability assumption which can lead to unintended consequences [guide2023fairness], e.g., disproportionally high wait times for ride-hails in remote areas [maciejewski2016assignment], discrimination against certain train types or routes in real-time train re-scheduling [luan2017non], or exacerbation of energy poverty [sousa2019peer].
Exact full comparability of the costs incurred by different agents can be impeded by various reasons:
  • Costs may be subjective and not straightforward to compare (e.g., perceived value of travel time, which depends on socio-economic and circumstantial factors[zamparini2007meta, liu2024adaptive]).
  • Costs may be difficult to quantify because they depend on complex models and unknown parameters (e.g., the yield of a farm as a function of the water allocated to it) or because they include information that is not available in the decision process.
  • The comparability framework is an elegant way to encode fairness considerations, i.e., what elements of agents’ costs are deemed relevant to the decision problem at hand (e.g., a deliberate decision not to account for traffic delays accumulated outside the road network of interest).
  • When costs are self-reported, they are vulnerable to strategic manipulation (e.g., customers declaring their electricity needs [muthirayan2019mechanism]).
@@
    State the level of interpersonal comparability. Such a decision is guided by how subjective agents’ preferences are and how well they can be measured, but it can also be a deliberate decision guided by politics, social agreements, and public sentiment.
@@
    Select the appropriate social cost function. Based on the determined comparability level, the choice of SCF is restricted to a specific functional form.
@@
    Align the control design to the social cost function. The design of the sociotechnical system must be guided uniquely by the explicit optimization of the selected SCF. Criteria that cannot be expressed through the SCF (e.g., rights, fairness constraints, external moral values, distributional constraints) should not be employed.
@@
We urge control engineers to adopt this rigorous and explainable process in all application domains where the design is guided by social cost/welfare functions [Munoz2023]. The mathematical guarantees that we provide can be incorporated in the design of system-wide metrics that are now being used in multi-agent control, like Price of Anarchy [Chandan2024, Zhang2018, Piliouras2017, Wang2017, hill2023tradeoff]. They can also guide the design of optimization dynamics [jadbabaie2009distributed, Menon2014, romvary2021proximal, Camisa2020OptNash] and resource allocation mechanisms [Farhadi2019, Maheswaran2004]. Finally, this work responds to the growing interest in fairness in control [CSSroadmap, jalota2021efficiency, villa2023fair, bang2024mobility, elokda2024carma, khargonekar2024climate, elokda2025vision], as many fairness notions arise naturally from careful considerations of the level of interpersonal comparability [elokda2025vision]. This approach can also clarify the implicit fairness assumptions in existing game-theoretic solution concepts, such as the variational Generalized Nash Equilibrium, which has been shown to rely on a strong, and often unstated, comparability assumption that, if violated, can lead to undesirable “fairness” outcomes [hall2025limits].
@@
The aforementioned three-step procedure can also be interpreted in reverse: if a specific notion of efficiency/fairness is desired (utilitarian, min/max, etc.) then a sufficient level of comparability needs to be achieved, which may require the collection of additional information from the agents, the design of a manipulation-safe mechanism, or the use of more accurate cost models. This may give a principled perspective on how freely agents’ costs can be designed and engineered [Marden2013, Jensen2018, Marden2014].
The third step of the procedure consists of using the selected SCF for control design. These SCFs are functionally similar to some objectives used in multi-objective optimization [Liu2001MOOOptandControl, Chinchuluun2007MOOsurvey], and specifically in multi-objective control design problems [Gambier2007MOOoverview, Gambier2008MPCandPIDMOO]. For example, linear-quadratic regulators can be designed to minimize weighted-sum [DeNicolao1993UtopianLQ], minmax [Li1990MinmaxLQ], and proportional [Shtessel1996proportionalLQR] costs. Similar cost functions can be incorporated in receding horizon control decisions [DeVito2007Recedinghorizon, Bemporad2009MultiobjectiveMPC, Zavala2012Utopiatracking] and in computational control design tools [Reynoso2017controllerTuning]. While these works do not address the aggregation of costs across multiple agents, they offer methods and algorithms for implementing social choice in practice.
The remainder of this paper is structured as follows. In Section 2, we set the stage for a principled welfarist approach based on social choice theory. In Section 3 we show how preferences can be properly aggregated into a SCF based on their level of interpersonal comparability. In Section 4 we illustrate the proposed design process via three examples of sociotechnical systems. Section 5 concludes the paper.
2 Welfarism
@@
In this section we introduce preliminaries adopted from social choice theory [DAspremont1977, roberts_interpersonal_1980, dAspremontGevers2002] that allow a rigorous aggregation of individual costs and derivation of social cost functions. We lay out the foundations of a so called “welfarist” approach, that requires that all the relevant information for the social decision is contained in the agents’ cost functions.
2.A Preliminaries and Axiomatic Foundations
@@
Let $\mathcal{N}=\{1,2,\dots,n\}$ be a finite set of agents and let $\mathcal{X}$ denote the set of feasible outcomes $x$. An outcome can be a specific allocation of a scarce resource or any control decision that affects the agents. Each agent $i$ attaches a (possibly negative) cost $J_{i}:\mathcal{X}\to\mathbb{R}$ to every outcome, with $J_{i}(\cdot)\in\mathcal{J}$, where $\mathcal{J}$ is the set of all real-valued functions on $\mathcal{X}$. We let $\mathbf{J}=(J_{1},\dots,J_{n})$ denote a profile of such costs for the entire set of agents. When formulating a resource allocation problem that depends on agents’ individual evaluations represented by $\mathbf{J}$, one needs to define a single social preference relation $\succsim_{\mathbf{J}}$ on $\mathcal{X}$, so as to capture the collective or “social” viewpoint. We first define a mapping between cost profiles of the agents and the resulting social preference, which can be expressed using Social Cost Functionals (SCFL):
Definition 1.
Let $\mathcal{R}$ be the set of all complete and transitive binary relations on $\mathcal{X}$. A Social Cost Functional (SCFL) is a mapping
$\mathfrak{F}\;:\;\mathcal{J}^{n}\;\longrightarrow\;\mathcal{R},$
that assigns a social preference relation $\succsim_{\mathbf{J}}=\mathfrak{F}(\mathbf{J})$ on $\mathcal{X}$ to each profile $\mathbf{J}=(J_{1},\dots,J_{n})\in\mathcal{J}^{n}$.
@@
Before determining how to incorporate each agent’s evaluation into a collective decision, one should first specify which fundamental properties (or axioms) this social preference relation $\succsim_{\mathbf{J}}$ needs to satisfy. The following two classical well-established properties can be expected from such a social preference relation.
@@
In control applications, it is convenient to work with real-valued measures of social cost rather than a social ordering of outcomes. This requires an additional mild continuity assumption, which postulates that if an outcome $x$ is socially strictly preferred over $y$, it should remain so under a small enough perturbation of the individual costs. As we show in Section 3.B, this condition is easily verified for the setting considered in this work.
Definition 2 (Pairwise Continuity (PC) [hammond2023roberts]).
Let $\mathbb{R}_{++}$ denote the set of strictly positive real numbers. For every $\varepsilon\in\mathbb{R}_{++}^{N}$, there exists $\varepsilon^{\prime}\in\mathbb{R}_{++}^{N}$ such that for every profile $\mathbf{J}$ and every pair $x,y\in\mathcal{X}$ with $x\,\succ_{\mathbf{J}}\,y$, there exists a profile $\tilde{\mathbf{J}}$ satisfying $\tilde{\mathbf{J}}(x)\geq\mathbf{J}(x)+\varepsilon^{\prime}$ and $\tilde{\mathbf{J}}(y)\leq\mathbf{J}(y)+\varepsilon$, and such that $x\,\succ_{\tilde{\mathbf{J}}}\,y.$
In some cases it may be desirable to enforce an additional axiom that prohibits distinguishing individual agents.
@@
2.B Welfarism
If we accept these axioms (P) and (IIA) as reasonable and impose them on our social preference relation, together with condition (PC), we can establish a welfarist principle [Sen1979_welfarism, DAspremont1977], that implies that any complete and transitive social ordering can be expressed through a Social Cost Function (SCF) $C$ that depends only on the preference profile $\mathbf{J}$.
Proposition 1 (Welfarism; Thm 3.7, [dAspremontGevers2002], Thm 1, [roberts_interpersonal_1980]).
@@ 2.B Welfarism
The SCF $C$ thus represents the SCFL $\mathfrak{F}$, and all relevant information for ranking outcomes is contained in the vector $\mathbf{J}(x)=\bigl(J_{1}(x),\dots,J_{n}(x)\bigr)$ of agents’ costs for each outcome $x$. The welfarist approach prescribes that the social choice must be guided exclusively by the social cost function $C(\mathbf{J}(x))$, rather than by any other characteristics of the decision $x$.
@@
As mentioned in the introduction, several well-known examples of SCFs $C$ have been proposed to aggregate individual costs. For instance, the utilitarian approach, often attributed to Bentham [bentham_introduction_1789] as later formalized by Harsanyi [harsanyi_cardinal_1955], aggregates costs via a sum: $C_{\mathrm{util}}(J_{1},\dots,J_{n})=\sum_{i=1}^{n}J_{i}$. By contrast, the Rawlsian or min-max rule, inspired by Rawls [rawls_theory_1971], focuses on the well-being of the worst-off agent: $C_{\mathrm{Rawls}}=\max_{i\in\mathcal{N}}J_{i}$. Another example is the Nash Social Welfare function [nash_bargaining_1950, kaneko1979nash], which is defined as $C_{\mathrm{Nash}}(J_{1},\dots,J_{n})=-\prod_{i=1}^{n}(-J_{i})$. Notably, these three canonical SCFs can all be viewed as members of the family of Hölder means. Generally, a large body of literature guides the selection of a SCF based on desirable notions of fairness, equity (e.g. envy-freeness and its relaxations, proportional fairness, Pigou-Dalton principle etc.), efficiency, or some trade-off between the two, reflected in measures such as maximin share (MMS) guarantee, alpha-fairness, or statistical inequality indices (e.g., the Gini coefficient, coefficient of variation, or Hoover index) [moulin_fair_2003, bertimas2011priceoffairness, guide2023fairness, bogomolnaia_competitive_2017]. Instead, the focus of this work is to provide a “first-principles” way to make this choice by first considering the fundamental assumptions of interpersonal comparability of the agents’ costs, which are needed to properly define “efficiency” and “fairness” in the first place, and from which the appropriate aggregation rule naturally follows.
3 Cost Aggregation and Comparability
@@
The welfarist principle (Proposition 1) tells us that any social ordering under (P), (IIA) and (PC) must be representable by a SCF that depends solely on the costs incurred by the agents. One must still decide which $C$ is appropriate in practice. Such a choice needs to be guided by a deliberate decision on how individual costs can be measured and compared across agents, so that meaningful interpersonal trade-offs can be made in the social choice. Addressing the interpersonal comparability issue corresponds to deciding which transformations of the individual costs leave the social ranking unchanged [DAspremont1977, roberts_interpersonal_1980]. By restricting or expanding the class of invariance transformations of the individual costs, one effectively selects a level of comparability for the agents’ costs. Once the comparability assumptions are fixed, the shape of the SCF $C$ is essentially pinned down.
3.A Comparability Levels
Definition 3 (Invariance Transformations under a SCFL).